OpenStudy (anonymous):

If \[\huge x=\sqrt[3]{7+5\sqrt{2}}-\frac{ 1 }{ \huge \sqrt[3]{7+5\sqrt{2}} }\] then find the value of \[\huge x ^{3}+3x-14\]

OpenStudy (anonymous):

@ganeshie8 @zepdrix

ganeshie8 (ganeshie8):

say \[\large x = t - \dfrac{1}{t}\] \[\begin{align} \\ x^3 + 3x-14 &= \left(t-\dfrac{1}{t}\right)^3 - 3 \left(t-\dfrac{1}{t}\right)-14\\~\\&=t^3 - \dfrac{1}{t^3} -3(t - \dfrac{1}{t}) +3(t - \dfrac{1}{t})-14 \\~\\ &=t^3 - \dfrac{1}{t^3}-14 \end{align} \]

ganeshie8 (ganeshie8):

plugin \(t^3 = 7 +5\sqrt{2} \) above ^^

OpenStudy (anonymous):

oh i got it

ganeshie8 (ganeshie8):

there os a simpler way

OpenStudy (anonymous):

how?

ganeshie8 (ganeshie8):

nvm, it wont work >.<

OpenStudy (anonymous):

what was the idea , it's ok

ganeshie8 (ganeshie8):

i think you should get 0 or 1 in the end as final answer

OpenStudy (anonymous):

perfect :)

ganeshie8 (ganeshie8):

corrected typo say \[\large x = t - \dfrac{1}{t}\] \[\begin{align} \\ x^3 \color{red}{+} 3x-14 &= \left(t-\dfrac{1}{t}\right)^3 \color{red}{+} 3 \left(t-\dfrac{1}{t}\right)-14\\~\\&=t^3 - \dfrac{1}{t^3} -3(t - \dfrac{1}{t}) +3(t - \dfrac{1}{t})-14 \\~\\ &=t^3 - \dfrac{1}{t^3}-14 \end{align} \]

OpenStudy (anonymous):

thanks